Mrs.Santos invests P21,000 at 8% interest rate.How much more must he invest at 9 1/2% interest to have an annual income of 9% of his total investments? (answer is P42,000 show your solution)
Answer by richwmiller(17219)

0.095*x+0.08*21000=0.09(21000+x)
0.095*x+1680=1890+0.09x
0.095x-0.09x=1890-1680
0.005x=210
x=42000.0 at 9.5% (3990.0) is added to 21000 at 8% (1680) to earn 9% of 63000.0 (5670.0)
check
0.095*42000.0+0.08*21000=0.09(21000+42000.0)
3990.0+1680=0.09(63000.0)
5670.0=5670.0
ok